OpenStudy (anonymous):

find the equation of the line with zero slope and passes through points ( 8, -3 )

OpenStudy (luigi0210):

Welcome to Openstudy @MathHelp1010 :)

OpenStudy (anonymous):

Thanks

OpenStudy (luigi0210):

Okay, first, do you know the point slope formula?

OpenStudy (anonymous):

Y=mx+b right ?

OpenStudy (luigi0210):

That's the slope formula :) Point slope is: y-y1=m(x-x1)

OpenStudy (anonymous):

OK

OpenStudy (luigi0210):

So all you have to do really is just plug in the values.

OpenStudy (anonymous):

y+3=0(x-8) Like that ?

OpenStudy (anonymous):

\[y+3=0(x-8)\]

OpenStudy (anonymous):

\[Y=3\]

OpenStudy (anonymous):

Oh I get what I did wrong , thanks

OpenStudy (loser66):

but to me, there is something wrong. It doesn't make sense at aaaalll.

OpenStudy (anonymous):

Oh

OpenStudy (loser66):

where Luigi? @Luigi0210

OpenStudy (anonymous):

They just stopped talking to me after giving me the equation

OpenStudy (loser66):

oh, I'm wrong, they are right. yes, it 's y = -3

OpenStudy (loser66):

so sorry friend

OpenStudy (luigi0210):

What?

OpenStudy (anonymous):

You are totally fine ! Thanks for checking :)

OpenStudy (loser66):